# Generalization Bounds via Meta-Learned Model Representations: PAC-Bayes and Sample Compression Hypernetworks

### v2

@@ B.2 Second General PAC-Bayes Sample Compression Theorem
$=\log\frac{1}{P_{J}(\operatorname{\mathbf{j}}^{*})}+\mathrm{KL}(Q_{\Omega}||P_{\Omega})$
@@
In the setting of Theorem 2.4, with $\Delta_{C}(q,p)=-\ln(1-(1-e^{-C})p)-Cq$ (where $C>0$), with probability at least $1-\delta$ over the draw of $S\sim\operatorname{\mathcal{D}}^{m}$, we have:
@@
$\operatorname*{\mathbb{E}}_{\omega\sim Q_{\Omega}}\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}(\operatorname{\mathscr{R}}(S_{\mathbf{j}},\omega))\leq\frac{1}{1-e^{-C}}\quantity[1-\exp\quantity(-C\operatorname*{\mathbb{E}}_{\omega\sim Q_{\Omega}}\widehat{\operatorname{\mathcal{L}}}_{S_{\overline{\operatorname{\mathbf{j}}}}}(\operatorname{\mathscr{R}}(S_{\mathbf{j}},\omega))-\frac{\mathrm{KL}(Q_{\Omega}||P_{\Omega})-\ln\quantity({P_{J}(\operatorname{\mathbf{j}})\cdot\delta})}{m-\max_{\operatorname{\mathbf{j}}\in J}\operatorname{|\mathbf{j}|}})]$
@@
In the setting of Theorem 2.4, with $\Delta(q,p)=\mathrm{kl}(q,p)=q\cdot\ln\frac{q}{p}+(1-q)\cdot\ln\frac{1-q}{1-p}$, with probability at least $1-\delta$ over the draw of $S\sim\operatorname{\mathcal{D}}^{m}$, we have:
@@
$~{}~{}~{}~{}~{}\mathrm{kl}\quantity(\operatorname*{\mathbb{E}}_{\omega\sim Q_{\Omega}}\widehat{\operatorname{\mathcal{L}}}_{S_{\overline{\operatorname{\mathbf{j}}}}}(\operatorname{\mathscr{R}}(S_{\mathbf{j}},\omega)),\operatorname*{\mathbb{E}}_{\omega\sim Q_{\Omega}}\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}(\operatorname{\mathscr{R}}(S_{\mathbf{j}},\omega)))\leq\frac{1}{m-\max_{\operatorname{\mathbf{j}}\in J}\operatorname{|\mathbf{j}|}}\quantity[\mathrm{KL}(Q_{\Omega}||P_{\Omega})+\ln\quantity(\frac{\operatorname*{\mathbb{E}}_{\operatorname{\mathbf{j}}\sim P_{J}}2\sqrt{m-\operatorname{|\mathbf{j}|}}}{P_{J}(\operatorname{\mathbf{j}})\cdot\delta})]$
@@
In the setting of Theorem 2.4, for any $\lambda>0$, with $\Delta(q,p)=\lambda(p-q)$, with a $\varsigma^{2}$-sub-Gaussian loss function $\ell:\operatorname{\mathcal{Y}}\times\operatorname{\mathcal{Y}}\to\operatorname{\mathbb{R}}$, with probability at least $1-\delta$ over the draw of $S\sim\operatorname{\mathcal{D}}^{m}$, for all $\operatorname{\mathbf{j}}\in J$ and posterior probability distribution $Q_{\Omega}$, we have:
$\operatorname*{\mathbb{E}}_{\omega\sim Q_{\Omega}}\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}(\operatorname{\mathscr{R}}(S_{\mathbf{j}},\omega))\leq\operatorname*{\mathbb{E}}_{\omega\sim Q_{\Omega}}\widehat{\operatorname{\mathcal{L}}}_{S_{\overline{\operatorname{\mathbf{j}}}}}(\operatorname{\mathscr{R}}(S_{\mathbf{j}},\omega))+\frac{1}{\lambda(m-\max_{\operatorname{\mathbf{j}}\in J}\operatorname{|\mathbf{j}|})}\quantity[\mathrm{KL}(Q_{\Omega}||P_{\Omega})+\ln\frac{1}{P_{J}(\operatorname{\mathbf{j}})\cdot\delta}+\ln\operatorname*{\mathbb{E}}_{\operatorname{\mathbf{j}}\sim P_{J}}\exp\quantity(\frac{(n-\operatorname{|\mathbf{j}|})\lambda^{2}\varsigma^{2}}{2})]$
@@
Let $\mathcal{Z}=\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}}$. We consider a set of sample-compressed predictors that can be defined using a reconstruction function $\operatorname{\mathscr{R}}$, a compression set $\operatorname{\mathbf{j}}$ and a message $\omega$. Following from the PAC-Bayes Sample compression work of Laviolette & Marchand (2005), we define the distribution on $J\times\Omega$ instead of the hypothesis class. This is equivalent, as the predictors are sample-compressed. Then, it is equivalent to sample $h$ from a distribution on the hypothesis class and to sample $(\operatorname{\mathbf{j}},\omega)$ from a distribution on $J\times\Omega$ and set $h=\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}},\omega)$.
@@
$\phi(h,S)=\exp\quantity(\frac{\alpha-1}{\alpha}(m-\max_{\operatorname{\mathbf{j}}\in J}\operatorname{|\mathbf{j}|})\Delta\quantity(\operatorname{\widehat{\mathcal{L}}}_{S_{\overline{\operatorname{\mathbf{j}}}}}(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}},\omega)),\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}},\omega)))).$
@@
$\frac{\alpha}{\alpha-1}\ln\phi(h,S)$ $=\frac{\alpha}{\alpha-1}\ln\exp\quantity(\frac{\alpha-1}{\alpha}(m-\max_{\operatorname{\mathbf{j}}\in J}\operatorname{|\mathbf{j}|})\Delta\quantity(\operatorname{\widehat{\mathcal{L}}}_{S_{\overline{\operatorname{\mathbf{j}}}}}(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}},\omega)),\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}},\omega))))$
$=\frac{\alpha}{\alpha-1}\frac{\alpha-1}{\alpha}(m-\max_{\operatorname{\mathbf{j}}\in J}\operatorname{|\mathbf{j}|})\Delta\quantity(\operatorname{\widehat{\mathcal{L}}}_{S_{\overline{\operatorname{\mathbf{j}}}}}(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}},\omega)),\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}},\omega)))$
$=(m-\max_{\operatorname{\mathbf{j}}\in J}\operatorname{|\mathbf{j}|})\Delta\quantity(\operatorname{\widehat{\mathcal{L}}}_{S_{\overline{\operatorname{\mathbf{j}}}}}(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}},\omega)),\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}},\omega))).$
@@
$\leq\ln\quantity[\operatorname*{\mathbb{E}}_{\operatorname{\mathbf{j}}\sim P_{J}}\operatorname*{\mathbb{E}}_{\omega\sim P_{\Omega}}\operatorname*{\mathbb{E}}_{S^{\prime}_{\operatorname{\mathbf{j}}}\sim\operatorname{\mathcal{D}}^{\operatorname{|\mathbf{j}|}}}\operatorname*{\mathbb{E}}_{S^{\prime}_{\overline{\operatorname{\mathbf{j}}}}\sim\operatorname{\mathcal{D}}^{m-\operatorname{|\mathbf{j}|}}}\exp\quantity((m-\operatorname{|\mathbf{j}|})\Delta\quantity(\operatorname{\widehat{\mathcal{L}}}_{S^{\prime}_{\overline{\operatorname{\mathbf{j}}}}}(\operatorname{\mathscr{R}}(S^{\prime}_{\operatorname{\mathbf{j}}},\omega)),\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}(\operatorname{\mathscr{R}}(S^{\prime}_{\operatorname{\mathbf{j}}},\omega))))]$
@@
With these two derivations, we have a disintegrated PAC-Bayes Sample Compression bound, with a posterior distribution $Q_{S}=B(S,P)$. We need to reframe this to accommodate the fact that we want a single compression set and a message sampled from an uncountable set of messages.
To do so, we use a compression function ${\mathcal{C}}$, that takes as input a dataset and outputs a compression set $S_{\operatorname{\mathbf{j}}^{\star}}={\mathcal{C}}(S)$, and an encoder $\operatorname{\mathcal{E}}$, that takes a dataset and outputs a message $\omega^{\star}$, defined such that for any learning algorithm $A$, we have $A(S)=\operatorname{\mathscr{R}}({\mathcal{C}}(S),\operatorname{\mathcal{E}}(S))$. We define the posterior distribution on the compression set as a Dirac distribution such that the only compression set with a non-zero probability is $\operatorname{\mathbf{j}}^{\star}$. For the message, the posterior distribution $Q_{\Omega}^{S}$ on the message is a distribution with mean $\omega^{\star}$ (for example a Gaussian). Thus, our algorithm $\zeta(S,P_{\Omega})$ returns $\quantity(S_{\operatorname{\mathbf{j}}^{\star}},Q_{\Omega}^{S})$. The Rényi Divergence then becomes :
@@
$D_{\alpha}(Q||P)$ $=\frac{1}{\alpha-1}\ln\quantity[\operatorname*{\mathbb{E}}_{(\operatorname{\mathbf{i}},\omega)\sim P}\quantity[\frac{Q(\operatorname{\mathbf{i}},\omega)}{P(\operatorname{\mathbf{i}},\omega)}]^{\alpha}]$
@@
$=\frac{1}{\alpha-1}\ln\quantity[\operatorname*{\mathbb{E}}_{(\operatorname{\mathbf{i}},\omega)\sim P}\quantity[\frac{Q_{J}(\operatorname{\mathbf{i}})Q_{\Omega}(\omega)}{P_{J}(\operatorname{\mathbf{i}})P_{\Omega}(\omega)}]^{\alpha}]$
@@
$=\frac{1}{\alpha-1}\ln\quantity[\operatorname*{\mathbb{E}}_{(\operatorname{\mathbf{i}},\omega)\sim P}\quantity[\frac{Q_{J}(\operatorname{\mathbf{i}})}{P_{J}(\operatorname{\mathbf{i}})}]^{\alpha}\quantity[\frac{Q_{\Omega}(\omega)}{P_{\Omega}(\omega)}]^{\alpha}]$
@@
$=\frac{1}{\alpha-1}\ln\quantity[\operatorname*{\mathbb{E}}_{\omega\sim P_{\Omega}}\operatorname*{\mathbb{E}}_{\operatorname{\mathbf{i}}\sim P_{J}}\quantity[\frac{Q_{J}(\operatorname{\mathbf{i}})}{P_{J}(\operatorname{\mathbf{i}})}]^{\alpha}\quantity[\frac{Q_{\Omega}(\omega)}{P_{\Omega}(\omega)}]^{\alpha}]$
@@
$=\frac{1}{\alpha-1}\ln\quantity[\operatorname*{\mathbb{E}}_{\omega\sim P_{\Omega}}\quantity[\frac{Q_{\Omega}(\omega)}{P_{\Omega}(\omega)}]^{\alpha}\operatorname*{\mathbb{E}}_{\operatorname{\mathbf{i}}\sim P_{J}}\quantity[\frac{Q_{J}(\operatorname{\mathbf{i}})}{P_{J}(\operatorname{\mathbf{i}})}]^{\alpha}]$
@@
$=\frac{1}{\alpha-1}\ln\quantity[\operatorname*{\mathbb{E}}_{\omega\sim P_{\Omega}}\quantity[\frac{Q_{\Omega}(\omega)}{P_{\Omega}(\omega)}]^{\alpha}P_{J}(\operatorname{\mathbf{j}})\quantity[\frac{1}{P_{J}(\operatorname{\mathbf{j}})}]^{\alpha}]$
$=\frac{1}{\alpha-1}\ln\quantity[\operatorname*{\mathbb{E}}_{\omega\sim P_{\Omega}}\quantity[\frac{Q_{\Omega}(\omega)}{P_{\Omega}(\omega)}]^{\alpha}\quantity[\frac{1}{P_{J}(\operatorname{\mathbf{j}})}]^{\alpha-1}]$
$=\frac{1}{\alpha-1}\ln\quantity[\operatorname*{\mathbb{E}}_{\omega\sim P_{\Omega}}\quantity[\frac{Q_{\Omega}(\omega)}{P_{\Omega}(\omega)}]^{\alpha}]+\frac{1}{\alpha-1}\ln\quantity[\quantity[\frac{1}{P_{J}(\operatorname{\mathbf{j}})}]^{\alpha-1}]$
$=\frac{1}{\alpha-1}\ln\quantity[\operatorname*{\mathbb{E}}_{\omega\sim P_{\Omega}}\quantity[\frac{Q_{\Omega}(\omega)}{P_{\Omega}(\omega)}]^{\alpha}]+\ln\quantity[\frac{1}{P_{J}(\operatorname{\mathbf{j}})}]$
$=D_{\alpha}(Q_{\Omega}||P_{\Omega})+\ln\quantity[\frac{1}{P_{J}(\operatorname{\mathbf{j}})}]$
@@
$\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}},\omega))\leq\operatorname{\widehat{\mathcal{L}}}_{S_{\overline{\operatorname{\mathbf{j}}}}}\quantity(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}},\omega))+\frac{1}{\lambda(m-\max_{\operatorname{\mathbf{j}}\in J}\operatorname{|\mathbf{j}|})}\quantity[\frac{2\alpha-1}{\alpha-1}\ln\frac{2}{\delta}+\ln\frac{1}{P_{J}(\operatorname{\mathbf{j}})}+D_{\alpha}(Q_{\Omega}||P_{\Omega})+\ln\operatorname*{\mathbb{E}}_{\operatorname{\mathbf{j}}\sim P_{J}}\exp\quantity(\frac{(n-\operatorname{|\mathbf{j}|})\lambda^{2}\varsigma^{2}}{2})].$
Appendix G Algorithm and Architectures
The following pseudocode depicts our Sample Compression Hypernetworks approach, while Figure 8, Figure 9 and Figure 8 respectively depict our PAC-Bayes, sample compression and PAC-Bayes sample compression approaches.

### v3

@@ B.2 Second General PAC-Bayes Sample Compression Theorem
$=\operatorname*{\mathbb{E}}_{\omega\sim Q_{\Omega}}\quantity[Q_{J}(\operatorname{\mathbf{j}}^{*})\log\frac{Q_{J}(\operatorname{\mathbf{j}}^{*})}{P_{J}(\operatorname{\mathbf{j}}^{*})}+\sum_{\operatorname{\mathbf{j}}\neq\operatorname{\mathbf{j}}^{*}}Q_{J}(\operatorname{\mathbf{j}})\log\frac{Q_{J}(\operatorname{\mathbf{j}})}{P_{J}(\operatorname{\mathbf{j}})}]+\operatorname*{\mathbb{E}}_{\operatorname{\mathbf{j}}\sim Q_{J}}\operatorname*{\mathbb{E}}_{\omega\sim Q_{\Omega}}\log\frac{Q_{\Omega}(\omega)}{P_{\Omega}(\omega)}$
$=\operatorname*{\mathbb{E}}_{\omega\sim Q_{\Omega}}\quantity[1\cdot\log\frac{1}{P_{J}(\operatorname{\mathbf{j}}^{*})}+\sum_{\operatorname{\mathbf{j}}\neq\operatorname{\mathbf{j}}^{*}}0\cdot\log\frac{0}{P_{J}(\operatorname{\mathbf{j}})}]+\operatorname*{\mathbb{E}}_{\operatorname{\mathbf{j}}\sim Q_{J}}\operatorname*{\mathbb{E}}_{\omega\sim Q_{\Omega}}\log\frac{Q_{\Omega}(\omega)}{P_{\Omega}(\omega)}$
@@
$=\log\frac{1}{P_{J}(\operatorname{\mathbf{j}}^{*})}+\mathrm{KL}(Q_{\Omega}||P_{\Omega})\,.\qed$
@@
$\operatorname*{\mathbb{E}}_{\omega\sim Q_{\Omega}}\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}(\operatorname{\mathscr{R}}(S_{\mathbf{j}},\omega))\leq\frac{1}{1-e^{-C}}\quantity[1-\exp\quantity(-C\operatorname*{\mathbb{E}}_{\omega\sim Q_{\Omega}}\widehat{\operatorname{\mathcal{L}}}_{S_{\overline{\operatorname{\mathbf{j}}}}}(\operatorname{\mathscr{R}}(S_{\mathbf{j}},\omega))-\frac{\mathrm{KL}(Q_{\Omega}||P_{\Omega})-\ln\quantity({P_{J}(\operatorname{\mathbf{j}})\cdot\delta})}{m-\max_{\operatorname{\mathbf{j}}\in J}\operatorname{|\mathbf{j}|}})].$
@@
$~{}~{}~{}~{}~{}\mathrm{kl}\quantity(\operatorname*{\mathbb{E}}_{\omega\sim Q_{\Omega}}\widehat{\operatorname{\mathcal{L}}}_{S_{\overline{\operatorname{\mathbf{j}}}}}(\operatorname{\mathscr{R}}(S_{\mathbf{j}},\omega)),\operatorname*{\mathbb{E}}_{\omega\sim Q_{\Omega}}\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}(\operatorname{\mathscr{R}}(S_{\mathbf{j}},\omega)))\leq\frac{1}{m-\max_{\operatorname{\mathbf{j}}\in J}\operatorname{|\mathbf{j}|}}\quantity[\mathrm{KL}(Q_{\Omega}||P_{\Omega})+\ln\quantity(\frac{\operatorname*{\mathbb{E}}_{\operatorname{\mathbf{j}}\sim P_{J}}2\sqrt{m-\operatorname{|\mathbf{j}|}}}{P_{J}(\operatorname{\mathbf{j}})\cdot\delta})].$
@@
In the setting of Theorem 2.4, for any $\lambda>0$, with $\Delta(q,p)=\lambda(p-q)$, with a $\varsigma^{2}$-sub-Gaussian loss function $\ell:\operatorname{\mathcal{Y}}\times\operatorname{\mathcal{Y}}\to\operatorname{\mathbb{R}}$, with probability at least $1-\delta$ over the draw of $S\sim\operatorname{\mathcal{D}}^{m}$, for all $\operatorname{\mathbf{j}}\in J$ and posterior probability distribution $Q_{\Omega}$, we have
$\operatorname*{\mathbb{E}}_{\omega\sim Q_{\Omega}}\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}(\operatorname{\mathscr{R}}(S_{\mathbf{j}},\omega))\leq$ $\operatorname*{\mathbb{E}}_{\omega\sim Q_{\Omega}}\widehat{\operatorname{\mathcal{L}}}_{S_{\overline{\operatorname{\mathbf{j}}}}}(\operatorname{\mathscr{R}}(S_{\mathbf{j}},\omega))$
$+\frac{1}{\lambda(m-\max_{\operatorname{\mathbf{j}}\in J}\operatorname{|\mathbf{j}|})}\quantity[\mathrm{KL}(Q_{\Omega}||P_{\Omega})+\ln\frac{1}{P_{J}(\operatorname{\mathbf{j}})\cdot\delta}+\ln\operatorname*{\mathbb{E}}_{\operatorname{\mathbf{j}}\sim P_{J}}\exp\quantity(\frac{(n-\operatorname{|\mathbf{j}|})\lambda^{2}\varsigma^{2}}{2})].$
@@
Let $\mathcal{Z}=\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}}$. We consider a set of sample-compressed predictors that can be defined using a reconstruction function $\operatorname{\mathscr{R}}$, a compression set $\operatorname{\mathbf{j}}$ and a message $\omega$. Following from the PAC-Bayes Sample compression work of Laviolette & Marchand (2005), we define the distribution on $J\times\Omega$ instead of the hypothesis class; it is equivalent to sample $h$ from a distribution on the hypothesis class and to sample $(\operatorname{\mathbf{j}}^{\star},\omega^{\star})$ from a distribution on $J\times\Omega$ and set $h=\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}^{\star}},\omega^{\star})$.
@@
$\phi(h,S)=\exp\quantity(\frac{\alpha-1}{\alpha}(m-\max_{\operatorname{\mathbf{j}}\in J}\operatorname{|\mathbf{j}|})\Delta\quantity(\operatorname{\widehat{\mathcal{L}}}_{S_{\overline{\operatorname{\mathbf{j}}}^{\star}}}(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}^{\star}},\omega^{\star})),\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}^{\star}},\omega^{\star})))).$
@@
$\frac{\alpha}{\alpha-1}\ln\phi(h,S)$ $=\frac{\alpha}{\alpha-1}\ln\exp\quantity(\frac{\alpha-1}{\alpha}(m-\max_{\operatorname{\mathbf{j}}\in J}\operatorname{|\mathbf{j}|})\Delta\quantity(\operatorname{\widehat{\mathcal{L}}}_{S_{\overline{\operatorname{\mathbf{j}}}^{\star}}}(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}^{\star}},\omega^{\star})),\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}^{\star}},\omega^{\star}))))$
$=\frac{\alpha}{\alpha-1}\frac{\alpha-1}{\alpha}(m-\max_{\operatorname{\mathbf{j}}\in J}\operatorname{|\mathbf{j}|})\Delta\quantity(\operatorname{\widehat{\mathcal{L}}}_{S_{\overline{\operatorname{\mathbf{j}}}^{\star}}}(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}^{\star}},\omega^{\star})),\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}^{\star}},\omega^{\star})))$
$=(m-\max_{\operatorname{\mathbf{j}}\in J}\operatorname{|\mathbf{j}|})\Delta\quantity(\operatorname{\widehat{\mathcal{L}}}_{S_{\overline{\operatorname{\mathbf{j}}}^{\star}}}(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}^{\star}},\omega^{\star})),\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}^{\star}},\omega^{\star}))).$
@@
$\leq\ln\quantity[\operatorname*{\mathbb{E}}_{\operatorname{\mathbf{j}}\sim P_{J}}\operatorname*{\mathbb{E}}_{\omega\sim P_{\Omega}}\operatorname*{\mathbb{E}}_{S^{\prime}_{\operatorname{\mathbf{j}}}\sim\operatorname{\mathcal{D}}^{\operatorname{|\mathbf{j}|}}}\operatorname*{\mathbb{E}}_{S^{\prime}_{\overline{\operatorname{\mathbf{j}}}}\sim\operatorname{\mathcal{D}}^{m-\operatorname{|\mathbf{j}|}}}\exp\quantity((m-\operatorname{|\mathbf{j}|})\Delta\quantity(\operatorname{\widehat{\mathcal{L}}}_{S^{\prime}_{\overline{\operatorname{\mathbf{j}}}}}(\operatorname{\mathscr{R}}(S^{\prime}_{\operatorname{\mathbf{j}}},\omega)),\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}(\operatorname{\mathscr{R}}(S^{\prime}_{\operatorname{\mathbf{j}}},\omega))))].$
@@
With these two derivations, we have a disintegrated PAC-Bayes Sample Compression bound, with a posterior distribution $Q_{S}=\zeta(S,P)$. We need to reframe this to accommodate the fact that we want a single compression set and a message sampled from an uncountable set of messages.
To do so, we start by choosing $P=P_{J}\times P_{\Omega}$ and $Q_{S}=Q_{J}^{S}\times Q_{\Omega}^{S}$. Moreover, we choose $\zeta$ such that $Q_{J}^{S}$ is always a Dirac distribution such that the only compression set with a non-zero probability is $\operatorname{\mathbf{j}}^{\star}$. The Rényi Divergence then becomes:
$D_{\alpha}(Q||P)$ $=\frac{1}{\alpha-1}\ln\quantity[\operatorname*{\mathbb{E}}_{(\operatorname{\mathbf{j}},\omega)\sim P}\quantity[\frac{Q(\operatorname{\mathbf{j}},\omega)}{P(\operatorname{\mathbf{j}},\omega)}]^{\alpha}]$
$=\frac{1}{\alpha-1}\ln\quantity[\operatorname*{\mathbb{E}}_{(\operatorname{\mathbf{j}},\omega)\sim P}\quantity[\frac{Q_{J}(\operatorname{\mathbf{j}})Q_{\Omega}(\omega)}{P_{J}(\operatorname{\mathbf{j}})P_{\Omega}(\omega)}]^{\alpha}]$
$=\frac{1}{\alpha-1}\ln\quantity[\operatorname*{\mathbb{E}}_{(\operatorname{\mathbf{j}},\omega)\sim P}\quantity[\frac{Q_{J}(\operatorname{\mathbf{j}})}{P_{J}(\operatorname{\mathbf{j}})}]^{\alpha}\quantity[\frac{Q_{\Omega}(\omega)}{P_{\Omega}(\omega)}]^{\alpha}]$
$=\frac{1}{\alpha-1}\ln\quantity[\operatorname*{\mathbb{E}}_{\omega\sim P_{\Omega}}\operatorname*{\mathbb{E}}_{\operatorname{\mathbf{j}}\sim P_{J}}\quantity[\frac{Q_{J}(\operatorname{\mathbf{j}})}{P_{J}(\operatorname{\mathbf{j}})}]^{\alpha}\quantity[\frac{Q_{\Omega}(\omega)}{P_{\Omega}(\omega)}]^{\alpha}]$
$=\frac{1}{\alpha-1}\ln\quantity[\operatorname*{\mathbb{E}}_{\omega\sim P_{\Omega}}\quantity[\frac{Q_{\Omega}(\omega)}{P_{\Omega}(\omega)}]^{\alpha}\operatorname*{\mathbb{E}}_{\operatorname{\mathbf{j}}\sim P_{J}}\quantity[\frac{Q_{J}(\operatorname{\mathbf{j}})}{P_{J}(\operatorname{\mathbf{j}})}]^{\alpha}]$
@@
$=\frac{1}{\alpha-1}\ln\quantity[\operatorname*{\mathbb{E}}_{\omega\sim P_{\Omega}}\quantity[\frac{Q_{\Omega}(\omega)}{P_{\Omega}(\omega)}]^{\alpha}\quantity(P_{J}(\operatorname{\mathbf{j}}^{\star})\quantity[\frac{Q_{J}(\operatorname{\mathbf{j}}^{\star})}{P_{J}(\operatorname{\mathbf{j}}^{\star})}]^{\alpha}+\sum_{\operatorname{\mathbf{j}}\neq\operatorname{\mathbf{j}}^{\star}}P_{J}(\operatorname{\mathbf{j}})\quantity[\frac{Q_{J}(\operatorname{\mathbf{j}})}{P_{J}(\operatorname{\mathbf{j}})}]^{\alpha})]$
@@
$=\frac{1}{\alpha-1}\ln\quantity[\operatorname*{\mathbb{E}}_{\omega\sim P_{\Omega}}\quantity[\frac{Q_{\Omega}(\omega)}{P_{\Omega}(\omega)}]^{\alpha}\quantity(P_{J}(\operatorname{\mathbf{j}}^{\star})\quantity[\frac{1}{P_{J}(\operatorname{\mathbf{j}}^{\star})}]^{\alpha}+\sum_{\operatorname{\mathbf{j}}\neq\operatorname{\mathbf{j}}^{\star}}P_{J}(\operatorname{\mathbf{j}})\quantity[\frac{0}{P_{J}(\operatorname{\mathbf{j}})}]^{\alpha})]$
@@
$=\frac{1}{\alpha-1}\ln\quantity[\operatorname*{\mathbb{E}}_{\omega\sim P_{\Omega}}\quantity[\frac{Q_{\Omega}(\omega)}{P_{\Omega}(\omega)}]^{\alpha}P_{J}(\operatorname{\mathbf{j}}^{\star})\quantity[\frac{1}{P_{J}(\operatorname{\mathbf{j}}^{\star})}]^{\alpha}]$
@@
$=\frac{1}{\alpha-1}\ln\quantity[\operatorname*{\mathbb{E}}_{\omega\sim P_{\Omega}}\quantity[\frac{Q_{\Omega}(\omega)}{P_{\Omega}(\omega)}]^{\alpha}\quantity[\frac{1}{P_{J}(\operatorname{\mathbf{j}}^{\star})}]^{\alpha-1}]$
@@
$=\frac{1}{\alpha-1}\ln\quantity[\operatorname*{\mathbb{E}}_{\omega\sim P_{\Omega}}\quantity[\frac{Q_{\Omega}(\omega)}{P_{\Omega}(\omega)}]^{\alpha}]+\frac{1}{\alpha-1}\ln\quantity[\quantity[\frac{1}{P_{J}(\operatorname{\mathbf{j}}^{\star})}]^{\alpha-1}]$
@@
$=\frac{1}{\alpha-1}\ln\quantity[\operatorname*{\mathbb{E}}_{\omega\sim P_{\Omega}}\quantity[\frac{Q_{\Omega}(\omega)}{P_{\Omega}(\omega)}]^{\alpha}]+\ln\quantity[\frac{1}{P_{J}(\operatorname{\mathbf{j}}^{\star})}]$
$=D_{\alpha}(Q_{\Omega}||P_{\Omega})+\ln\quantity[\frac{1}{P_{J}(\operatorname{\mathbf{j}}^{\star})}]\,.\qed$
@@
$\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}},\omega))\leq$$\operatorname{\widehat{\mathcal{L}}}_{S_{\overline{\operatorname{\mathbf{j}}}}}\quantity(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{j}}},\omega))$
$+\frac{1}{\lambda(m-\max_{\operatorname{\mathbf{j}}\in J}\operatorname{|\mathbf{j}|})}\quantity[\frac{2\alpha-1}{\alpha-1}\ln\frac{2}{\delta}+\ln\frac{1}{P_{J}(\operatorname{\mathbf{j}})}+D_{\alpha}(Q_{\Omega}||P_{\Omega})+\ln\operatorname*{\mathbb{E}}_{\operatorname{\mathbf{j}}\sim P_{J}}\exp\quantity(\frac{(n-\operatorname{|\mathbf{j}|})\lambda^{2}\varsigma^{2}}{2})].$
Appendix G Algorithm
The following pseudocode depicts our Sample Compression Hypernetworks approach.